# Token-Specific Watermarking with Enhanced Detectability and Semantic Coherence for Large Language Models

### v2

@@ Abstract
Large language models generate high-quality responses with potential misinformation, underscoring the need for regulation by distinguishing AI-generated and human-written texts. Watermarking is pivotal in this context, which involves embedding hidden markers in texts during the LLM inference phase, which is imperceptible to humans. Current watermarking algorithms, however, face the challenge of achieving both the detectability of inserted watermarks and the semantic integrity of generated texts, where enhancing one aspect often undermines the other. To overcome this, we introduce a novel multi-objective optimization (MOO) approach for watermarking that utilizes lightweight networks to generate token-specific watermarking logits and splitting ratios. By leveraging MOO to optimize for both detection and semantic objective functions, our method simultaneously achieves detectability and semantic integrity. Experimental results show that our method outperforms current watermarking techniques in enhancing the detectability of texts generated by LLMs while maintaining their semantic coherence. Our code is available at https://github.com/mignonjia/TS_watermark.
Machine Learning, ICML
[…]
University of California, San Diego
[…]
{mhuo, ssomayaj, p1xie}@ucsd.edu
@@ 1 Introduction
Various classification-based approaches have been proposed to determine whether a text is generated by humans or LLMs (OpenAI, 2023; Guo et al., 2023; Brown et al., 1992). However, as LLM-generated texts increasingly match the quality of human-generated ones, these methods are losing their robustness. For instance, OpenAI released its own AI classifier (OpenAI, 2023) in early 2023, but it was later withdrawn due to its low accuracy. Recently, watermarking techniques in LLMs have gained popularity (Abdelnabi & Fritz, 2021; He et al., 2022a; Zhang et al., 2023; Kirchenbauer et al., 2023a; Kuditipudi et al., 2023; Aaronson, 2023; Liu et al., 2023; Ren et al., 2023; Lee et al., 2023; Wang et al., 2023; Wouters, 2023). These techniques embed hidden patterns in LLM-generated texts that, while imperceptible to humans, can be detected by algorithms. Kirchenbauer et al. (2023a) propose an effective watermarking method, referred to as KGW in this paper. To generate a new token with a watermark, the hash of the preceding token is used as a random seed to divide the vocabulary into a green list and a red list with a fixed splitting ratio (i.e., the percentage of tokens in the green list). Then a constant value, known as the watermark logit, is added to the LLM-produced logits on the green list tokens. This adjustment increases the probability of selecting the ‘green’ tokens as the new token. To identify the presence of a watermark, a statistical test is carried out. This involves counting the number of green list tokens in the generated text. A higher incidence of green tokens suggests that the text is likely to contain a watermark.
@@
The design of KGW emphasizes easy detection of watermarked texts. However, this approach often compromises the semantic coherence of the texts, as highlighted in Lee et al. (2023). One primary cause of this issue is that KGW uses a constant splitting ratio and watermark logit across all tokens, without taking into account the context and semantics of the specific token being generated. For instance, given the prefix “The sun rises in the”, it is highly probable that the next token to be generated should be “east”. But KGW’s watermarking mechanism might not choose “east” if the splitting ratio is low and the watermark logit is high. A lower splitting ratio results in a smaller number of words being added to the green list, thus significantly reducing the chances of including the word “east” in it. On the other hand, a higher watermark logit raises the probability of choosing a token from this list. Consequently, it becomes highly unlikely for “east” to be selected as the next token for generation, which will significantly impact the semantic coherence of the text. To mitigate this issue, adjusting the splitting ratio and watermark logit is necessary, perhaps by increasing the splitting ratio or decreasing the watermark logit. Unfortunately, KGW lacks such an adaptive mechanism. Conversely, if KGW utilizes a low, constant splitting ratio and a small, uniform watermark logit for every token to maintain semantic integrity, this approach may hinder the watermark’s detectability. Hence, it is crucial to adaptively assign token-specific values for these hyperparameters, simultaneously ensuring high detectability while maintaining semantic integrity. Some works have been proposed to improve the semantic coherence of watermarked texts. Liu et al. (2023) introduce a semantically invariant watermarking method that ensures a high correlation between the watermark logits and input text embeddings, consequently enhancing semantic performance. Kuditipudi et al. (2023) reduce the semantics distortion by ensuring an unbiased distribution of tokens before and after watermarking, using an exponential minimum sampling strategy which can positively influence semantics. However, these works face challenges in enhancing both detectability and semantic coherence at the same time: improving one aspect frequently compromises the other.
@@ 2.1 Watermarking Methods
Inference-time watermarking methods insert statistical signals into model logits during inference to improve detectability. KGW (Kirchenbauer et al., 2023a) comes under this umbrella. However, text watermarked by this method suffers from reduced semantic coherence. A series of works have been proposed to address this issue. Wang et al. (2023) introduce a constraint stipulating that the perplexity of watermarked text must not exceed the perplexity of the original text, as produced by the same language model without watermarking, by more than a small margin. While perplexity can serve as a measure of the quality of generated text, it fails to account for the semantic proximity between the watermarked text and its original version. This oversight can lead to semantic disparities between the two texts. Chen et al. (2023) propose to split the vocabulary into a green list and red list in a semantically more balanced manner so that any token in the red list can be replaced by a token in the green list. However, their splitting method is based on LLaMA2 (Touvron et al., 2023), which is computationally expensive. Furthermore, Liu et al. (2023) introduce a semantically invariant watermarking approach, based on the motivation that watermark logits should derive from the semantics of preceding tokens, not their token IDs. They employ Compositional-BERT (Chanchani & Huang, 2023) to extract semantic representations of preceding tokens and produce watermark logits from these representations. However, this method incurs high computation overhead during inference due to the utilization of Compositional-BERT. Wouters (2023) present a closed-form solution for optimizing watermarking logits to enhance semantic integrity, with a predetermined splitting ratio. However, this solution relies on assumptions that may not hold in real-world scenarios, such as all tokens following identical distribution, and the absence of distribution shifts post-watermarking. Additionally, their analysis does not extend to determining an optimal splitting ratio - a factor that is critical for balancing detectability and semantic preservation. To enhance semantic coherence of watermarked computer programs, SWEET (Lee et al., 2023) proposes to selectively insert watermarks into high-entropy tokens instead of every token. While this method is effective for watermarking code, it demonstrates weak detection capabilities in texts across general domains.
@@ 2.2 Multi-Objective Optimization
Multi-objective optimization (MOO) (Deb et al., 2016) addresses the challenge of simultaneously optimizing several objectives which often conflict with each other. The Multiple-gradient Descent Algorithm (MGDA) (Désidéri, 2012) is a notable gradient-based approach designed for solving MOO problems. MGDA aims to identify a single gradient direction theoretically capable of optimizing all objectives concurrently. If the gradients are not normalized, the direction of the optimization is expected to be mostly influenced by the gradients of small norms (Désidéri, 2012). To mitigate this, various normalization methods have been developed (Chen et al., 2018; Milojkovic et al., 2019), striving for a more equitable consideration of all involved gradients. In our work, we employ MGDA to identify Pareto optimal solutions between detectability and semantic coherence.
@@ 5 Experiments
In our experiments, we insert watermarks to two prominent LLMs: OPT-1.3B (Zhang et al., 2022) and LLAMA2-7B (Touvron et al., 2023). The evaluation of watermarking is based on three aspects: the trade-off between detectability and semantic integrity, robustness against attacks of the watermarks, and computational time complexity. We also perform an analysis of our learned $\delta$ and $\gamma$, and present an ablation study in Appendix H. Training details including the hyperparameter settings can be found in Appendix D.
@@ 5.1 Experimental Settings
Following Kirchenbauer et al. (2023a), we utilize texts from the news-like subset of the C4 dataset (Raffel et al., 2019) to insert watermarks. For each text from this dataset, the last 200 tokens are truncated and designated as the ‘baseline completion’ (i.e., human-written texts). The remaining tokens from the start of the text are considered the ‘prompt’. Conditioned on this prompt, the LLM generates a token sequence equivalent in length to the baseline completion, incorporating watermarks within this generation. Following Kirchenbauer et al. (2023a), we filter the dataset to include texts ranging from 500 to 1000 tokens. This dataset is then divided randomly into three subsets: 6,400 texts for training, 500 for validation, and another 500 for testing. In text generation, multinomial sampling with a temperature of $1.0$ is used.
@@
We compare our methods with KGW (Kirchenbauer et al., 2023a) and SWEET (Lee et al., 2023) (which is based on KGW), by plotting trade-off curves of TPR and semantic similarity for the two settings, FPR=0% and FPR=1%. For each setting, we set $\gamma=0.25$ and vary $\delta$ in $[1.0,3.0]$ for both KGW and SWEET to generate different pairs of TPR and semantic similarity. For our method, we generate different pairs of TPR and semantic similarity using different initializations of the $\gamma$- and $\delta$-generator networks. Then we fit a curve to these pairs (i.e., points) for each method and plot the curves to form the Pareto frontier (Giagkiozis & Fleming, 2014), which represents the best trade-off curve of TPR and semantic similarity. For more details, please refer to Appendix D.
@@ 5.2 Results on Detectability and Semantic Coherence
Figure 2 and 3 show results on OPT-1.3B and LLAMA2-7B, respectively, which highlight three key observations. First, our approach improves the Pareto frontier compared to the KGW baseline, attributable to two key factors. The first factor is that our method uniquely learns token-specific splitting ratios and watermark logits, which take into account the distinct context and semantics of each token. This is critical because the number of semantically appropriate tokens that can be chosen as the next token in the sequence can vary substantially at different time steps during text generation, depending on the context. In scenarios where this number is small (such as the example in Sec. 1), it is essential to lower the splitting ratio and watermark logit to reduce the watermark strength. This adjustment increases the likelihood of selecting these valid choices as the next token, thereby maintaining semantics. Conversely, in situations where this number is large, increasing the splitting ratio and watermark logit to an appropriate level can improve detectability while posing minimal risk to semantic coherence. Our method has the adaptability that allows for token-specific adjustments in splitting ratios and watermark logits, while KGW employs uniform values across all tokens. The second factor distinguishing our approach is the incorporation of multi-objective optimization, which enables simultaneous maximization of detectability and semantic integrity. This is achieved by concurrently optimizing a differentiable detection loss and a semantic loss. In contrast, the KGW method cannot explicitly optimize for these two objectives at the same time.
Second, our method outperforms the SWEET baseline with an improved Pareto frontier at an FPR 0%, and it demonstrates comparable performance at an FPR of 1%. Notably, at an FPR of 0%, SWEET is less effective compared to our approach and KGW. This could stem from its selective watermarking strategy, which was initially designed for code watermarking. SWEET selectively watermarks only high-entropy words, leaving low-entropy words unwatermarked ($\delta=0$). However, our experiments in the general domain, specifically with the C4 dataset, show that SWEET does not consistently achieve a 100% TPR even in regions of high $\delta$, implying a strong watermark. For example, at $(\gamma,\delta)=(0.25,3.0)$, upon examining the generated samples with z-scores below 4.0 (the z-score threshold at 0% FPR for SWEET), we observed that these samples contain, on average, only 7 high-entropy tokens eligible for SWEET watermarking. This number is significantly lower than the total generation size of 200 tokens, restricting SWEET’s detectability even with a high $\delta$ strength. Conversely, our method and KGW achieve a 100% TPR in high $\delta$ regions. Another limitation of SWEET is its reliance on the original prompts for detection, a requirement that is often impractical in real-world scenarios. Thus, we introduce ‘SWEET_no_prompt’, which evaluates entropy without taking prompts into account. ‘SWEET_no_prompt’ underperforms our method at an FPR of 0% and within high $\delta$ regions at an FPR of 1%, highlighting the method’s dependency on prompts, which is an impractical assumption.
Figure 3: Performance of our model (trained on OPT-1.3B) and KGW when applied to LLAMA2-7B.
@@
Third, our method demonstrates good generalization capacity across LLMs. As shown in Figure 3, our model ($\gamma$- and $\delta$-generator networks), initially trained on OPT-1.3B, demonstrates a better Pareto frontier when applied to LLAMA2-7B. This adaptability is likely because our method learns the watermarking parameters that reflect the general nature of language itself (see Sec. 5.3 and Appendix F for discussions), not just on the specific details of the LLMs.

### v3

@@ Abstract
Large language models generate high-quality responses with potential misinformation, underscoring the need for regulation by distinguishing AI-generated and human-written texts. Watermarking is pivotal in this context, which involves embedding hidden markers in texts during the LLM inference phase, which is imperceptible to humans. Achieving both the detectability of inserted watermarks and the semantic quality of generated texts is challenging. While current watermarking algorithms have made promising progress in this direction, there remains significant scope for improvement. To address these challenges, we introduce a novel multi-objective optimization (MOO) approach for watermarking that utilizes lightweight networks to generate token-specific watermarking logits and splitting ratios. By leveraging MOO to optimize for both detection and semantic objective functions, our method simultaneously achieves detectability and semantic integrity. Experimental results show that our method outperforms current watermarking techniques in enhancing the detectability of texts generated by LLMs while maintaining their semantic coherence. Our code is available at https://github.com/mignonjia/TS_watermark.
watermarking, large language models, text generation
@@ 1 Introduction
Various classification-based approaches have been proposed to determine whether a text is generated by humans or LLMs (OpenAI, 2023; Guo et al., 2023b; Brown et al., 1992). However, as LLM-generated texts increasingly match the quality of human-generated ones, these methods are losing their robustness. For instance, OpenAI released its own AI classifier (OpenAI, 2023) in early 2023, but it was later withdrawn due to its low accuracy. Recently, watermarking techniques in LLMs have gained popularity (Abdelnabi & Fritz, 2021; He et al., 2022a; Zhang et al., 2023; Kirchenbauer et al., 2023a; Kuditipudi et al., 2023; Aaronson, 2023; Liu et al., 2023; Ren et al., 2023; Lee et al., 2023; Wang et al., 2023; Wouters, 2023). These techniques embed hidden patterns in LLM-generated texts that, while imperceptible to humans, can be detected by algorithms. Kirchenbauer et al. (2023a) propose an effective watermarking method, referred to as KGW in this paper. To generate a new token with a watermark, the hash of the preceding token is used as a random seed to divide the vocabulary into a green list and a red list with a fixed splitting ratio (i.e., the percentage of tokens in the green list). Then a constant value, known as the watermark logit, is added to the LLM-produced logits on the green list tokens. This adjustment increases the probability of selecting the ‘green’ tokens as the new token. To identify the presence of a watermark, a statistical test is carried out. This involves counting the number of green list tokens in the generated text. A higher incidence of green tokens suggests that the text is likely to contain a watermark.
@@
The design of KGW emphasizes easy detection of watermarked texts. However, this approach often compromises the semantic coherence of the texts, as highlighted in Lee et al. (2023). One primary cause of this issue is that KGW uses a constant splitting ratio and watermark logit across all tokens, without taking into account the context and semantics of the specific token being generated. For instance, given the prefix “The sun rises in the”, it is highly probable that the next token to be generated should be “east”. But KGW’s watermarking mechanism might not choose “east” if the splitting ratio is low and the watermark logit is high. A lower splitting ratio results in a smaller number of words being added to the green list, thus significantly reducing the chances of including the word “east” in it. On the other hand, a higher watermark logit raises the probability of choosing a token from this list. Consequently, it becomes highly unlikely for “east” to be selected as the next token for generation, which will significantly impact the semantic coherence of the text. To mitigate this issue, adjusting the splitting ratio and watermark logit is necessary, perhaps by increasing the splitting ratio or decreasing the watermark logit. Unfortunately, KGW lacks such an adaptive mechanism. Conversely, if KGW utilizes a low, constant splitting ratio and a small, uniform watermark logit for every token to maintain semantic integrity, this approach may hinder the watermark’s detectability. Hence, it is crucial to adaptively assign token-specific values for these hyperparameters, simultaneously ensuring high detectability while maintaining semantic integrity. Some works have been proposed to improve the semantics of watermarked texts. Lee et al. (2023) proposes a selective watermarking strategy to preserve semantics by adding watermarks only to high-entropy tokens while preserving the original logits for low-entropy tokens. Kuditipudi et al. (2023) reduce the semantics distortion by ensuring an unbiased distribution of tokens before and after watermarking, using an exponential minimum sampling strategy which can positively influence semantics. However, these works face challenges in enhancing both detectability and semantic coherence at the same time: improving one aspect frequently compromises the other.
@@ 2.1 Watermarking Methods
Inference-time watermarking methods insert statistical signals into model logits during inference to improve detectability. KGW (Kirchenbauer et al., 2023a) comes under this umbrella. However, text watermarked by this method suffers from reduced semantic coherence. A series of works have been proposed to address this issue. Wang et al. (2023) introduce a constraint stipulating that the perplexity of watermarked text must not exceed the perplexity of the original text, as produced by the same language model without watermarking, by more than a small margin. However, this strategy is developed through a series of approximations to ensure practical applicability. These approximations can lead to semantic disparities between the two texts. Chen et al. (2023) propose to split the vocabulary into a green list and red list in a semantically more balanced manner so that any token in the red list can be replaced by a token in the green list. However, their splitting method is based on LLAMA2 (Touvron et al., 2023), which is computationally expensive. Furthermore, Liu et al. (2023) introduce a semantically invariant watermarking approach, aimed to improve attack and security robustness. They employ Compositional-BERT (Chanchani & Huang, 2023) to extract semantic representations of preceding tokens and produce watermark logits from these representations. However, this method incurs high computation overhead during inference due to the utilization of Compositional-BERT. Wouters (2023) present a closed-form solution for optimizing watermarking logits to enhance semantic integrity, with a predetermined splitting ratio. However, this solution relies on assumptions that may not hold in real-world scenarios, such as all tokens following identical distribution, and the absence of distribution shifts post-watermarking. Additionally, their analysis does not extend to determining an optimal splitting ratio - a factor that is critical for balancing detectability and semantic preservation. To enhance semantic coherence of watermarked computer programs, SWEET (Lee et al., 2023) proposes to selectively insert watermarks into high-entropy tokens instead of every token. While this method is effective for watermarking code, it demonstrates weak detection capabilities in texts across general domains.
@@ 2.2 Multi-Objective Optimization
Multi-objective optimization (MOO) (Deb et al., 2016) addresses the challenge of simultaneously optimizing several objectives which often conflict with each other. The Multiple-gradient Descent Algorithm (MGDA) (Désidéri, 2012) is a notable gradient-based approach designed for solving MOO problems. MGDA aims to identify a single gradient direction theoretically capable of optimizing all objectives concurrently. If the gradients are not normalized, the direction of the optimization is expected to be mostly influenced by the gradients of small norms (Désidéri, 2012). To mitigate this, various normalization methods have been developed (Chen et al., 2018; Milojkovic et al., 2019), striving for a more equitable consideration of all involved gradients. In our work, we employ MGDA to identify the Pareto optimal solutions between detectability and semantic coherence.
@@ 5 Experiments
In our experiments, we insert watermarks to two prominent LLMs: OPT-1.3B (Zhang et al., 2022), LLAMA2-7B, 13B and 70B (Touvron et al., 2023). The evaluation of watermarking is based on three aspects: the trade-off between detectability and semantic integrity, computational complexity, and robustness against attacks of the watermarks. We also perform an analysis of our learned $\delta$ and $\gamma$, and present an ablation study in Appendix K. Training details including the hyperparameter settings can be found in Appendix D.
@@ 5.1 Experimental Settings
Following Kirchenbauer et al. (2023a), we utilize texts from the news-like subset of the C4 dataset (Raffel et al., 2019) to insert watermarks. For each text from this dataset, the last 200 tokens are truncated and designated as the ‘baseline completion’ (i.e., human-written texts). The remaining tokens from the start of the text are considered the ‘prompt’. Conditioned on this prompt, the LLM generates a token sequence equivalent in length to the baseline completion, incorporating watermarks within this generation. Following Kirchenbauer et al. (2023a), we filter the dataset to include texts ranging from 500 to 1000 tokens. This dataset is then divided randomly into three subsets: 6,400 texts for training, 500 for validation, and another 500 for testing.
@@
We compare our methods against KGW (Kirchenbauer et al., 2023a), SWEET (Lee et al., 2023), MultiBit (Wang et al., 2023), SIR (Liu et al., 2023), and EXP-edit (Kuditipudi et al., 2023). Both SWEET and SIR rely on prompts during detection, which is impractical in many real-world scenarios. To ensure a fair comparison, we also include variants of these baselines that do not use prompts during detection, namely $\text{SWEET}_{\text{NoPrompt}}$ and $\text{SIR}_{\text{NoPrompt}}$. We plot trade-off curves of TPR and semantic similarity for the two settings, FPR=0% and FPR=1%. We vary $\delta$ within the appropriate range for each baseline to generate different pairs of TPR and semantic similarity. For KGW, SWEET, and $\text{SWEET}_{\text{NoPrompt}}$, we set $\gamma=0.25$. For MultiBit, we use message lengths of 9 and 7 for which the FPR is 0% and 1%, respectively. Multinomial sampling with a temperature of $1.0$ and Top-$k$ of 50 is used for text generation. For our method, we generate different pairs of TPR and semantic similarity using different initializations of the $\gamma$- and $\delta$-generator networks. Then we fit a curve to these pairs (i.e., points) for each method and plot the curves to form the Pareto frontier (Giagkiozis & Fleming, 2014), which represents the best trade-off curve of TPR and semantic similarity. For more details, please refer to Appendix D.
@@ 5.2 Results on Detectability and Semantic Coherence
Figure 2 and 3 show results on OPT-1.3B and LLAMA2 (7B, 13B and 70B), respectively, which highlight several key observations when comparing with the baselines.
Comparison with KGW (Kirchenbauer et al., 2023a): Our approach improves the Pareto frontier compared to the KGW baseline, attributable to two key factors. The first factor is that our method uniquely learns token-specific splitting ratios and watermark logits, which take into account the distinct context and semantics of each token. This is critical because the number of semantically appropriate tokens that can be chosen as the next token in the sequence can vary substantially at different time steps during text generation, depending on the context. In scenarios where this number is small (such as the example in Sec. 1), it is essential to lower the splitting ratio and watermark logit to reduce the watermark strength. This adjustment increases the likelihood of selecting these valid choices as the next token, thereby maintaining semantics. Conversely, in situations where this number is large, increasing the splitting ratio and watermark logit to an appropriate level can improve detectability while posing minimal risk to semantic coherence. Our method has the adaptability that allows for token-specific adjustments in splitting ratios and watermark logits, while KGW employs uniform values across all tokens. The second factor distinguishing our approach is the incorporation of multi-objective optimization, which enables simultaneous maximization of detectability and semantic integrity. This is achieved by concurrently optimizing a differentiable detection loss and a semantic loss. In contrast, the KGW method cannot explicitly optimize for these two objectives together.
@@
Comparison with SWEET (Lee et al., 2023): Our method achieves a better Pareto frontier than SWEET at 0% FPR while maintaining similar performance at 1% FPR. At 0% FPR, SWEET is notably less effective compared to our approach and KGW, and does not achieve 100% TPR. This may be due to its selective watermarking strategy, which targets only high-entropy words and leaves low-entropy words un-watermarked ($\delta=0$). For instance, at $(\gamma,\delta)=(0.25,3.0)$, an analysis of LLM-generated texts that SWEET failed to detect at 0% FPR shows that, on average, only 7 out of 200 tokens are high-entropy and eligible for SWEET watermarking. This limited number of watermarkable tokens reduces SWEET’s detectability even with high $\delta$ (see Appendix G). In contrast, both our method and KGW achieve a 100% TPR in high $\delta$ regions. Furthermore, $\text{SWEET}_{\text{NoPrompt}}$ underperforms SWEET, indicating the method’s dependence on prompts, which is impractical.
Comparison with MultiBit (Wang et al., 2023): Our method achieves a superior Pareto frontier compared to MultiBit. MultiBit embeds multi-bit information into LLM-generated texts, detecting a watermark if the decoded message matches the embedded one. However, embedding multi-bit information often reduces text quality. To mitigate this, Balance-Marking strategy is introduced to decrease the perplexity of watermarked texts. However, this method is developed through a series of approximations to ensure practical applicability, which might limit its effectiveness. In contrast, our method directly maximizes differentiable metrics of semantic coherence and detectability through multi-objective optimization, inherently improving them.
Comparison with SIR (Liu et al., 2023): Our method achieves an improved Pareto frontier than SIR. SIR primarily aims to enhance attack and security robustness, lacking a direct approach to boost text quality. Conversely, our method employs multi-objective optimization to effectively enhance both text quality and watermark detectability simultaneously. Additionally, $\text{SIR}_{\text{NoPrompt}}$ significantly underperforms compared to SIR, indicating a strong dependence on prompts while detection. SIR is less robust than SWEET in the no-prompt scenario, as it exhibits a greater performance degradation without prompts compared to SWEET.
Comparison with EXP-edit (Kuditipudi et al., 2023): EXP-edit’s generation process, which employs exponential minimum sampling, is pseudo-random and becomes deterministic with a watermark key, facilitating detection. Unlike ours and other KGW-based methods, which shift the distribution using $\delta$, EXP-edit does not alter the output distribution of LLMs, making it indistinguishable. Since EXP-edit does not have a $\delta$ parameter to vary and plot the Pareto frontier, we present the results in Table 1. Our method outperforms EXP-edit in both detectability and SimCSE, demonstrating that learning token-specific parameters to watermark enables appropriately shifting the output distribution to enhance detectability without significantly affecting semantics. This offers more freedom to effectively embed watermark compared to EXP-edit, which lacks this capability (Piet et al., 2023).
Table 1: Comparison of EXP-edit and Our Method
Method TPR @ 0% TPR @ 1% SimCSE
EXP-edit 0.922 0.996 0.655
EXP-edit (Top-$k$=50) 0.968 0.996 0.677
Ours (Top-$k$=50) 1.000 1.000 0.713
Since our decoding strategy uses the default Top-$k$ value of 50, we modified the original EXP-edit generation process, which did not implement Top-$k$, to sample from the top 50 tokens, naming it EXP-edit (Top-$k$=50). We observe that our method still outperforms EXP-edit (Top-$k$=50) in terms of detectability and SimCSE. Moreover, indistinguishable methods like EXP-edit, which rely on pseudo-random sampling during generation, cannot easily extend to other decoding methods like beam search, which do not involve randomness. In contrast, KGW-based methods can be applied on top of any decoding method such as beam search (Kirchenbauer et al., 2023a). See Appendix H for further discussion.
Computation Overhead: We evaluate the computational time of our method and the baselines for text generation and detection in Table 2. Our method achieves higher speeds than EXP-edit, SIR, and MultiBit, while achieving speeds comparable to KGW, SWEET, and No Watermarking.
Table 2: Generation and detection speed on OPT-1.3B for generating 200 tokens, measured in seconds. We also show memory utilization in Table 5.
Method Generation (s) Detection (s)
No Watermark 3.220 -
KGW 3.827 0.067
SWEET 4.030 0.127
EXP-edit 24.693 155.045
SIR 8.420 0.337
MultiBit 6.500 0.610
Ours 3.946 0.166
(a) LLAMA2-7B
(b) LLAMA2-13B
(c) LLAMA2-70B
Figure 3: Performance of Ours (trained on OPT-1.3B) and KGW when applied to LLAMA2 7B, 13B, and 70B.
@@
Generalizability: Our method demonstrates good generalizability across LLMs of different sizes. As shown in Figure 3, our model ($\gamma$- and $\delta$-generator networks), initially trained on OPT-1.3B, demonstrates a better Pareto frontier when applied to LLAMA2 7B, 13B and 70B. This adaptability is likely because our method learns the watermarking parameters that reflect the general nature of language itself (see Sec. 5.3 and Appendix I for discussions), not just on the specific details of the LLMs.